OpenStudy (anonymous):

Solve for n. (4 7/12+8 2/3)+6 5/12 = n+11 i will medal and fan

OpenStudy (anonymous):

nikeboi101

OpenStudy (gabbyalicorn):

4 7/12+8 2/3+6 5/12=n+11 Simplify both sides of the equation. 593=n+11 Flip the equation. n+11=593 Subtract 11 from both sides. n+11−11=593−11 n=263 Answer n=823

OpenStudy (gabbyalicorn):

8 2/3*

OpenStudy (anonymous):

thank you so much!!!!!

OpenStudy (anonymous):

mathmale i will fan you since you are going to help

OpenStudy (mathmale):

although y our equation, (4 7/12+8 2/3)+6 5/12 = n+11 , can be solved as is, you might find it easier to visualize if you'd (please) write it like this:\[\frac{ 47 }{ 12 }+\frac{ 82 }{ 3 }+\frac{ 65 }{ 12 }=n+11\]

OpenStudy (mathmale):

first, find the LCD and apply it, so that ALL terms have the same denominator.